# Deep Submillimeter and Radio Observations in the SSA22 Field. III. Multiwavelength Identifications and Properties of the 850 -selected Submillimeter Galaxies

### v1

@@ Abstract
We present a multiwavelength analysis of 850 -selected submillimeter galaxies (SMGs; deblended S 1 mJy) in the SSA22 field, where our deepest JCMT/SCUBA-2 observations reach a sensitivity of 0.80 mJy beam-1. Using multiple identification methods, we have identified 248 deblended SMG candidates for 192 SCUBA-2 sources. The average multiplicity of SCUBA-2 sources is 26%, with brighter sources exhibiting higher multiplicity. After applying quality cuts based on spectral energy distribution (SED) fitting reliability, our final sample comprises 221 SMGs associated with 186 SCUBA-2 sources. The SSA22 SMGs have a median infrared luminosity of (2.25 0.25) 1012 L⊙, with 63% ( 8%) of the sample classified as ULIRGs (HLIRGs). The median redshift of the sample is , while optically faint galaxies exhibit higher median redshift (). The comoving volume density of SMGs increases by a factor of at , plateauing at 1.78-3.16 10-5 cMpc-3 over 1-3 (including the overdensity). The significant overdensity of SMGs within large-scale structures demonstrates their reliability as tracers of cosmic structure formation at high redshift. The median stellar mass and SFR of our SMG sample are M⊙ and M⊙ yr-1, respectively. We observe a clear “downsizing” signature: after cosmic noon (), massive SMGs exhaust their gas reservoirs and transition to quiescence, while lower-mass SMGs continue forming stars and dominate the cosmic SFR density. The sample has a median dust mass of (1.95 0.14) 109 M⊙. The dust fraction (), which has a median value of (1.4 0.2) 10-2, is also influenced by “downsizing” effect and shows significant evolution at . The median V-band attenuation of SMGs is 3.09 0.07 mag. Although shows no strong evolution over , a combination of our data with literature measurements indicates a significant drop in dust attenuation at , suggesting a transition to less obscured star formation, possibly due to insufficient metal enrichment in the early universe.
@@ I Introduction
Facilities like Spitzer, Herschel, IRAM 30-m, ASTE, APEX, and SPT have greatly contributed, though their coarse angular resolution (Hodge and da Cunha, 2020) has been a limitation. Interferometers like SMA, NOEMA, and especially ALMA have revolutionized the field with high resolution and sensitivity, enabling detailed studies of individual high-redshift SMGs, including their gas dynamics (Hodge and da Cunha, 2020). Umehata et al. (2017, 2018); Franco et al. (2018); Hatsukade et al. (2018); Gómez-Guijarro et al. (2022) mapped GOODS-S and SSA22 in a wide region utilizing ALMA, while the Ex-MORA survey has assembled the largest homogeneous ALMA continuum dataset to date in the COSMOS field (Casey et al., 2021; Long et al., 2024). Typical ALMA surveys achieve CIB resolution fractions of 40–50% (Chen et al., 2022), and—again leveraging gravitational lensing—the ALCS survey successfully resolved 80% of the CIB (Fujimoto et al., 2024). The UDS and COSMOS fields host the largest samples of ALMA follow-up observations (Liu et al., 2019; Adscheid et al., 2024). Collectively, these efforts have expanded SMG studies in depth, breadth, and methodological sophistication.
@@
These observations have revealed that SMGs are a heterogeneous population, exhibiting significant scatter relative to the star-forming main sequence (Michałowski et al., 2017; Shim et al., 2022), yet they are consistently massive, dust-obscured star-forming galaxies. 850 -selected SMGs typically possess stellar masses of M⊙, with a median near M⊙, and reside in massive dark matter halos ( M⊙; Chen et al., 2016; An et al., 2019; Lim et al., 2020b). Their star formation rates (SFRs) range from tens to thousands of solar masses per year. Sustaining such an active process requires substantial gas reservoirs, with typical SMGs hosting gas masses of M⊙ and gas fractions 0.5-0.9 (miettinen2017; birkin2021; liao2024), fueled either by major mergers or environmental accretion. Their high redshift, dust obscuration, and intense star formation result in extremely red colors and high infrared luminosities L⊙, with cold dust temperatures of 20-40 K (swinbank2014; Dudzevičiūtė et al., 2020b). SMGs also show elevated X-ray detection rates. The literature shows active galactic nucleus (AGN) are common in SMGs, with incidence rates of 20–40% (Alexander et al., 2005a, 2008; Pope et al., 2008; Laird et al., 2010; Georgantopoulos et al., 2010; Johnson et al., 2013; Wang et al., 2013; Umehata et al., 2015), and brighter SMGs are more likely to host AGN (Wang et al., 2013; Cowie et al., 2018; Lim et al., 2020b). In some cases, SMGs are also identified as optically-selected quasars, where concurrent black hole growth and star formation drive rapid galaxy evolution (Alexander et al., 2005a, b). These characteristics strongly suggest that SMGs are progenitors of present-day massive elliptical galaxies, which hosting supermassive black holes (SMBHs) at their centers (Genzel et al., 2003; simpson2017; Lim et al., 2020a).
@@
SMGs, as massive star-forming galaxies, are frequently located in overdense environments and trace large-scale structures (Menéndez-Delmestre et al., 2013; Dannerbauer et al., 2014; Arrigoni Battaia et al., 2018; Zhang et al., 2024), in line with the general tendency of massive galaxies to reside in dense regions (Dressler, 1980). Umehata et al. (2015) first demonstrated—free from source confusion—that SMGs anchor galaxy overdensities within cosmic web filaments, and Umehata et al. (2019) later revealed the co-evolution of galaxies, gas, and black holes within these cosmic structures. Thus, SMGs are valuable tracers of overdense regions (Tamura et al., 2009; Umehata et al., 2019), though their utility may decline below due to downsizing and decreasing baryon conversion efficiency (Chapman et al., 2009; Miller et al., 2015, 2018), and they may incompletely trace the most massive overdensities (Miller et al., 2015).
@@
The formation mechanisms of SMGs are debated. Early views, influenced by their resemblance to scaled-up local ULIRGs, favored major gas-rich mergers (swinbank2014; Hyun et al., 2023), supported by morphological, kinematic, and simulation studies (Toft et al., 2014). Alternatively, some argue that mergers may be secondary, with SMGs spending significant time in a steady state of evolution (Daddi et al., 2007). Many SMGs lie on the high-mass end of the star-forming main sequence (Michałowski et al., 2012; cunha2015; Dunlop et al., 2017; Zavala et al., 2018; Lim et al., 2020b). Recent ALMA and JWST studies find that of SMGs possess undisturbed extended disks (Gillman et al., 2023, 2024; Le Bail et al., 2024; Chan et al., 2025; Hodge et al., 2025; Umehata et al., 2025; Zhang et al., 2025), suggesting that internal instabilities or minor perturbations can trigger intense star formation and dust production in isolated disks (Gillman et al., 2024). Numerical models also support sustained starbursts in gas-rich disks (swinbank2014).
@@
The S2CLS is a major JCMT/SCUBA-2 program that mapped seven extragalactic deep fields, covering a total area 5 deg2 and a depth of mJy (Geach et al., 2017). SCUBA-2 sources in fields such as COSMOS (Michałowski et al., 2017; An et al., 2019), UDS (Michałowski et al., 2017), EGS (Zavala et al., 2018), and NEP (Shim et al., 2022) have been extensively analyzed. As one of the S2CLS target fields, SSA22 field was well observed; however, to better exploit its potential for studying SMGs, we obtained deeper SCUBA-2 integrations toward SSA22 and complemented them with coordinated VLA 3 GHz observations (Ao et al., 2017; Zeng et al., 2024). The SSA22 field has been extensively studied over the past two decades across various wavelengths and cosmic structures (C. C. Steidel, K. L. Adelberger, M. Dickinson, M. Giavalisco, M. Pettini, and M. Kellogg (1998); 107; Y. Matsuda, T. Yamada, T. Hayashino, H. Tamura, R. Yamauchi, T. Murayama, T. Nagao, K. Ohta, S. Okamura, M. Ouchi, K. Shimasaku, Y. Shioya, and Y. Taniguchi (2005); B. D. Lehmer, D. M. Alexander, J. E. Geach, I. Smail, A. Basu-Zych, F. E. Bauer, S. C. Chapman, Y. Matsuda, C. A. Scharf, M. Volonteri, and T. Yamada (2009b); B. D. Lehmer, D. M. Alexander, S. C. Chapman, I. Smail, F. E. Bauer, W. N. Brandt, J. E. Geach, Y. Matsuda, J. R. Mullaney, and A. M. Swinbank (2009a); T. M. A. Webb, T. Yamada, J.-S. Huang, M. L. N. Ashby, Y. Matsuda, E. Egami, M. Gonzalez, and T. Hayashimo (2009); M. Kubo, Y. K. Uchimoto, T. Yamada, M. Kajisawa, T. Ichikawa, Y. Matsuda, M. Akiyama, T. Hayashino, M. Konishi, T. Nishimura, K. Omata, R. Suzuki, I. Tanaka, T. Yoshikawa, D. M. Alexander, G. G. Fazio, J.-S. Huang, and B. D. Lehmer (2013); M. Kubo, T. Yamada, T. Ichikawa, M. Kajisawa, Y. Matsuda, I. Tanaka, and H. Umehata (2015); H. Umehata, Y. Tamura, K. Kohno, B. Hatsukade, K. S. Scott, M. Kubo, T. Yamada, R. J. Ivison, R. Cybulski, I. Aretxaga, J. Austermann, D. H. Hughes, H. Ezawa, T. Hayashino, S. Ikarashi, D. Iono, R. Kawabe, Y. Matsuda, H. Matsuo, K. Nakanishi, T. Oshima, T. Perera, T. Takata, G. W. Wilson, and M. S. Yun (2014); H. Umehata, M. Fumagalli, I. Smail, Y. Matsuda, A. M. Swinbank, S. Cantalupo, C. Sykes, R. J. Ivison, C. C. Steidel, A. E. Shapley, J. Vernet, T. Yamada, Y. Tamura, M. Kubo, K. Nakanishi, M. Kajisawa, B. Hatsukade, and K. Kohno (2019); H. Umehata, M. Kubo, I. Smail, B. D. Lehmer, E. B. Monson, K. Nakanishi, and Y. Matsuda (2025); Y. Kato, Y. Matsuda, I. Smail, A. M. Swinbank, B. Hatsukade, H. Umehata, I. Tanaka, T. Saito, D. Iono, Y. Tamura, K. Kohno, D. K. Erb, B. D. Lehmer, J. E. Geach, C. C. Steidel, D. M. Alexander, T. Yamada, and T. Hayashino (2016); Y. Ao, Y. Matsuda, C. Henkel, D. Iono, D. M. Alexander, S. C. Chapman, J. Geach, B. Hatsukade, M. Hayes, N. K. Hine, Y. Kato, R. Kawabe, K. Kohno, M. Kubo, M. Lehnert, M. Malkan, K. M. Menten, T. Nagao, R. P. Norris, M. Ouchi, T. Saito, Y. Tamura, Y. Taniguchi, H. Umehata, and A. Weiss (2017); O. R. Cooper, C. M. Casey, J. A. Zavala, J. B. Champagne, E. da Cunha, A. S. Long, J. S. Spilker, and J. Staguhn (2022); X. Zeng, Y. Ao, and Y. Zhang (2024); S. Huang, H. Umehata, I. Smail, K. Nakanishi, B. Hatsukade, M. Kubo, Y. Tamura, T. Saito, and S. Ikarashi (2025)), amassing rich multiwavelength data from X-ray to radio. It offers a unique laboratory for probing both early galaxy evolution and the assembly of cosmic web structures.
@@
In this paper, we combine 850 and 3 GHz data with extensive multi-wavelength ancillary datasets to identify SMG counterparts, perform spectral energy distribution (SED) fitting, and investigate their physical properties. A key challenge is the coarse resolution of single-dish submillimeter telescopes like SCUBA-2, requiring counterpart identification at higher-resolution. Although bands with positive -corrections and varying depths can miss higher-redshift or warmer SMGs (swinbank2014), they successfully identify the majority within the expected redshift range. Radio and 24 identifications are highly reliable and widely used (Chen et al., 2016), although these methods still miss a fraction of real SMGs (Hodge et al., 2013). The ADF22 survey (Umehata et al., 2017, 2018) provides additional counterparts for our sample. To enhance the completeness and maximize the potential of multiwavelength data, IRAC 8 sources have been widely utilized (Ashby et al., 2006; Michałowski et al., 2012, 2017; Zavala et al., 2018). Numerous studies employ various optical/near-infrared (NIR) wavelengths to identify high-redshift extremely red galaxies (e.g., HIEROs: Wang et al. (2012); KIEROs: Wang et al. (2016); OIRTC: Chen et al. (2016)).
@@ II.1 SCUBA-2 850
We utilize a stacked point-spread function (PSF) and a “top-down” iterative source extraction algorithm. Compared to the S2CLS SSA22 map, our deeper SCUBA-2 map reveals 212 new sources, of which 145 lie in the central region of the field (where 1.2 mJy) and 122 have flux densities greater than 4 mJy. While most of these newly detected sources are at low signal-to-noise ratios (SNR ), we still recover 93 sources at significance. Accounting for the false detection rate (Zeng et al., 2024), we estimate that approximately 170 of these new detections are authentic submillimeter sources.

### v2

@@ Abstract
We present a multiwavelength analysis of 850 -selected submillimeter galaxies (SMGs; deblended S 1 mJy) in the SSA22 field, using JCMT/SCUBA-2 data reaching 0.80 mJy beam-1. We identify 248 SMG candidates for 192 SCUBA-2 sources, with a multiplicity rate of 26% that increases with flux. After quality cuts based on SED fitting reliability, our final sample contains 221 SMGs associated with 186 SCUBA-2 sources. The SSA22 SMGs have a median infrared luminosity of (2.25 0.25) 1012 L⊙, with 63% ( 8%) of the sample classified as ULIRGs (HLIRGs). The median redshift is (higher, , for optically faint galaxies). SMGs trace large-scale structure, showing a significant overdensity at . The comoving SMG number density rises by a factor of at , plateauing at cMpc-3 over 1-3 (including the overdensity). The sample has median stellar mass M⊙, SFR M⊙ yr-1, dust mass M⊙, and mag. We detect a clear “downsizing” signature: , massive SMGs quench, while lower-mass systems dominate later star formation. The dust-to-stellar mass ratio, median , evolves significantly below . Although remains stable over , combined literature data suggest a sharp drop at , indicating a transition to less obscured star formation likely due to low metal enrichment in the early universe.
@@ I Introduction
Facilities like Spitzer, Herschel, IRAM 30-m, ASTE, APEX, and SPT have greatly contributed, though their coarse angular resolution (Hodge and da Cunha, 2020) has been a limitation. Interferometers like SMA, NOEMA, and especially ALMA have revolutionized the field with high resolution and sensitivity, enabling detailed studies of individual high-redshift SMGs, including their gas dynamics (Hodge and da Cunha, 2020). Umehata et al. (2017, 2018); Franco et al. (2018); Hatsukade et al. (2018); Gómez-Guijarro et al. (2022) mapped GOODS-S and SSA22 in a wide region utilizing ALMA, while the Ex-MORA survey has assembled the largest homogeneous ALMA continuum dataset to date in the COSMOS field (Casey et al., 2021; Long et al., 2024). Typical ALMA surveys achieve CIB resolution fractions of 40–50% (Chen et al., 2022), and—again leveraging gravitational lensing—the ALCS survey successfully resolved 80% of the CIB (Fujimoto et al., 2024). The UDS and COSMOS fields host the largest samples of ALMA follow-up observations (Stach et al., 2019; Dudzevičiūtė et al., 2020b; simpson2020; Liu et al., 2019; Adscheid et al., 2024). Collectively, these efforts have expanded SMG studies in depth, breadth, and methodological sophistication.
@@
These observations have revealed that SMGs are a heterogeneous population, exhibiting significant scatter relative to the star-forming main sequence (Shim et al., 2022), yet they are consistently massive, dust-obscured star-forming galaxies. 850 -selected SMGs typically possess stellar masses of M⊙, with a median near M⊙, and reside in massive dark matter halos ( M⊙; Lim et al., 2020b; Stach et al., 2021). Their star formation rates (SFRs) range from tens to thousands of solar masses per year. Sustaining such an active process requires substantial gas reservoirs, with typical SMGs hosting gas masses of M⊙ and gas fractions 0.5-0.9 (miettinen2017; birkin2021; liao2024), fueled either by major mergers or environmental accretion. Their high redshift, dust obscuration, and intense star formation result in extremely red colors and high infrared luminosities L⊙, with cold dust temperatures of 20-40 K (swinbank2014; Dudzevičiūtė et al., 2020b). The fraction of SMGs hosting active galactic nuclei (AGNs) has been a subject of debate in earlier studies. Initial work suggested AGN fractions as high as 20–40% (Alexander et al., 2005a; Pope et al., 2008; Laird et al., 2010; Georgantopoulos et al., 2010; Johnson et al., 2013; Wang et al., 2013). However, studies in the ALMA era have revealed that these earlier estimates were likely inflated. More recent analyses indicate that the AGN fraction among SMGs is below 20%, typically in the range of 15-20% (Wang et al., 2013; An et al., 2019; Stach et al., 2019; uematsu2025). Moreover, brighter and more massive SMGs appear more likely to host AGNs (Wang et al., 2013; Cowie et al., 2018; Lim et al., 2020b). SMGs residing in high-redshift large-scale structures may also exhibit elevated AGN fractions (Umehata et al., 2015). Additionally, SMGs hosting AGNs show a higher incidence of major mergers (uematsu2025), suggesting that major mergers play a key role in triggering black hole activity in these systems. In some cases, SMGs are also identified as optically-selected quasars, where concurrent black hole growth and star formation drive rapid galaxy evolution (Alexander et al., 2005a, b). These characteristics strongly suggest that SMGs are progenitors of present-day massive elliptical galaxies, which hosting supermassive black holes (SMBHs) at their centers (Genzel et al., 2003; simpson2017; Lim et al., 2020a).
@@
Observations suggest that SMGs are often associated with overdense structures at high redshift ( 1.5 ) (Menéndez-Delmestre et al., 2013; Dannerbauer et al., 2014; Arrigoni Battaia et al., 2018; Smail, 2024; Zhang et al., 2024), and broadly consistent with the general tendency for massive galaxies to reside in denser environments (Dressler, 1980). Umehata et al. (2015) first demonstrated—free from source confusion—that SMGs anchor galaxy overdensities within cosmic web filaments, and Umehata et al. (2019) later revealed the co-evolution of galaxies, gas, and black holes within these cosmic structures. Thus, SMGs have been considered potential tracers of overdense regions (Tamura et al., 2009; Umehata et al., 2019), though their utility may decline below due to “downsizing” and decreasing baryon conversion efficiency (Chapman et al., 2009; Miller et al., 2015, 2018), and they may incompletely trace the most massive overdensities (Miller et al., 2015).
@@
The formation mechanisms of SMGs are debated. Early views, influenced by their resemblance to scaled-up local ULIRGs, favored major gas-rich mergers (swinbank2014; Hyun et al., 2023), supported by morphological, kinematic, and simulation studies (Toft et al., 2014). Alternatively, some argue that mergers may be secondary (Daddi et al., 2007). Many SMGs lie on the high-mass end of the star-forming main sequence (cunha2015; Dunlop et al., 2017; Zavala et al., 2018; Lim et al., 2020b). Recent ALMA and JWST studies find that of SMGs possess undisturbed extended disks (Gillman et al., 2023, 2024; Le Bail et al., 2024; Chan et al., 2025; Hodge et al., 2025; Umehata et al., 2025a; Zhang et al., 2025), suggesting that internal instabilities or minor perturbations can trigger intense star formation and dust production in isolated disks (Gillman et al., 2024). Numerical models also support sustained starbursts in gas-rich disks (swinbank2014).
@@
The S2CLS is a major JCMT/SCUBA-2 program that mapped seven extragalactic deep fields, covering a total area 5 deg2 and a depth of mJy (Geach et al., 2017). SCUBA-2 sources in fields such as COSMOS (An et al., 2019; simpson2019; simpson2020), UDS (Stach et al., 2018, 2019; Dudzevičiūtė et al., 2020b), EGS (Zavala et al., 2018), and NEP (Shim et al., 2022) have been extensively analyzed. As one of the S2CLS target fields, SSA22 field was well observed; however, to better exploit its potential for studying SMGs, we obtained deeper SCUBA-2 integrations toward SSA22 and complemented them with coordinated VLA 3 GHz observations (Ao et al., 2017; Zeng et al., 2024). The SSA22 field has been extensively studied over the past two decades across various wavelengths and cosmic structures (C. C. Steidel, K. L. Adelberger, M. Dickinson, M. Giavalisco, M. Pettini, and M. Kellogg (1998); 113; Y. Matsuda, T. Yamada, T. Hayashino, H. Tamura, R. Yamauchi, T. Murayama, T. Nagao, K. Ohta, S. Okamura, M. Ouchi, K. Shimasaku, Y. Shioya, and Y. Taniguchi (2005); B. D. Lehmer, D. M. Alexander, J. E. Geach, I. Smail, A. Basu-Zych, F. E. Bauer, S. C. Chapman, Y. Matsuda, C. A. Scharf, M. Volonteri, and T. Yamada (2009b); B. D. Lehmer, D. M. Alexander, S. C. Chapman, I. Smail, F. E. Bauer, W. N. Brandt, J. E. Geach, Y. Matsuda, J. R. Mullaney, and A. M. Swinbank (2009a); T. M. A. Webb, T. Yamada, J.-S. Huang, M. L. N. Ashby, Y. Matsuda, E. Egami, M. Gonzalez, and T. Hayashimo (2009); M. Kubo, Y. K. Uchimoto, T. Yamada, M. Kajisawa, T. Ichikawa, Y. Matsuda, M. Akiyama, T. Hayashino, M. Konishi, T. Nishimura, K. Omata, R. Suzuki, I. Tanaka, T. Yoshikawa, D. M. Alexander, G. G. Fazio, J.-S. Huang, and B. D. Lehmer (2013); M. Kubo, T. Yamada, T. Ichikawa, M. Kajisawa, Y. Matsuda, I. Tanaka, and H. Umehata (2015); H. Umehata, Y. Tamura, K. Kohno, B. Hatsukade, K. S. Scott, M. Kubo, T. Yamada, R. J. Ivison, R. Cybulski, I. Aretxaga, J. Austermann, D. H. Hughes, H. Ezawa, T. Hayashino, S. Ikarashi, D. Iono, R. Kawabe, Y. Matsuda, H. Matsuo, K. Nakanishi, T. Oshima, T. Perera, T. Takata, G. W. Wilson, and M. S. Yun (2014); H. Umehata, M. Fumagalli, I. Smail, Y. Matsuda, A. M. Swinbank, S. Cantalupo, C. Sykes, R. J. Ivison, C. C. Steidel, A. E. Shapley, J. Vernet, T. Yamada, Y. Tamura, M. Kubo, K. Nakanishi, M. Kajisawa, B. Hatsukade, and K. Kohno (2019); H. Umehata, M. Kubo, I. Smail, B. D. Lehmer, E. B. Monson, K. Nakanishi, and Y. Matsuda (2025a); Y. Kato, Y. Matsuda, I. Smail, A. M. Swinbank, B. Hatsukade, H. Umehata, I. Tanaka, T. Saito, D. Iono, Y. Tamura, K. Kohno, D. K. Erb, B. D. Lehmer, J. E. Geach, C. C. Steidel, D. M. Alexander, T. Yamada, and T. Hayashino (2016); Y. Ao, Y. Matsuda, C. Henkel, D. Iono, D. M. Alexander, S. C. Chapman, J. Geach, B. Hatsukade, M. Hayes, N. K. Hine, Y. Kato, R. Kawabe, K. Kohno, M. Kubo, M. Lehnert, M. Malkan, K. M. Menten, T. Nagao, R. P. Norris, M. Ouchi, T. Saito, Y. Tamura, Y. Taniguchi, H. Umehata, and A. Weiss (2017); O. R. Cooper, C. M. Casey, J. A. Zavala, J. B. Champagne, E. da Cunha, A. S. Long, J. S. Spilker, and J. Staguhn (2022); X. Zeng, Y. Ao, and Y. Zhang (2024); S. Huang, H. Umehata, I. Smail, K. Nakanishi, B. Hatsukade, M. Kubo, Y. Tamura, T. Saito, and S. Ikarashi (2025)), amassing rich multiwavelength data from X-ray to radio. It offers a unique laboratory for probing both early galaxy evolution and the assembly of cosmic web structures.
@@
In this paper, we combine 850 and 3 GHz data with extensive multi-wavelength ancillary datasets to identify SMG counterparts, perform spectral energy distribution (SED) fitting, and investigate their physical properties. A key challenge is the coarse resolution of single-dish submillimeter telescopes like SCUBA-2, requiring counterpart identification at higher-resolution. Although bands with positive -corrections and varying depths can miss higher-redshift or warmer SMGs (swinbank2014), they successfully identify the majority within the expected redshift range. Radio and 24 identifications are reliable and widely used (Chen et al., 2016), although these methods still miss a fraction of real SMGs (Hodge et al., 2013; An et al., 2018). The ADF22 survey (Umehata et al., 2017, 2018) provides additional counterparts for our sample. To enhance the completeness and maximize the potential of multiwavelength data, IRAC 8 sources have also been widely utilized (Ashby et al., 2006; Zavala et al., 2018). Numerous studies employ various optical/near-infrared (NIR) wavelengths (Smail et al., 1999, 2002; Frayer et al., 2004) to identify high-redshift extremely red galaxies (e.g., HIEROs: Wang et al. (2012); KIEROs: Wang et al. (2016); OIRTC: Chen et al. (2016)).
@@ II.1 SCUBA-2 850
We utilize a stacked point-spread function (PSF) and a “top-down” iterative source extraction algorithm. Jackknife maps are constructed, and Monte Carlo simulations are employed to estimate the completeness and false detection rates. Compared to the S2CLS SSA22 map, our deeper SCUBA-2 map reveals 212 new sources, of which 145 lie in the central region of the field (where 1.2 mJy) and 122 have flux densities greater than 4 mJy. While most of these newly detected sources are at low signal-to-noise ratios (SNR ), we still recover 93 sources at significance. Accounting for the false detection rate (Zeng et al., 2024), we estimate that approximately 170 of these new detections are authentic submillimeter sources.